OpenStudy (anonymous):

does anyone know how to do this one. write the difference in expanded form 893- 514 using exponents

OpenStudy (anonymous):

not sure how to do this

OpenStudy (jdoe0001):

well, I'm not sure what they mean... but it sounds like 893 - 514 = 379 now we can say that we can EXPAND 379 to be a sum like 81+36+25 + 169 + 49 + 19

OpenStudy (jdoe0001):

\(\bf 81+36+25 + 169 + 49 + 19 \implies 9^2+6^2+5^2+13^2+7^2+19\)

OpenStudy (anonymous):

I think she wants me to write it something like this https://angel.spcollege.edu/AngelUploads/QuestionData/37d2f850-c4fd-4c48-9749-8fc615336511/36232G36223323412361.png# {0af42309-0f91-4d18-921b-c02083637aef}

OpenStudy (anonymous):

need to write it like that and dont remember how

OpenStudy (jdoe0001):

ok, so we know that the difference is 379 so can you provide.... say 3 numbers, that added together will give 379? 1st number must end with "00", 2nd number must end with "0" and the last one will be what's remaining of the 379

OpenStudy (anonymous):

lost

OpenStudy (jdoe0001):

can you think of a number that ends with "00" <--- zero zero

OpenStudy (jdoe0001):

lemm show you what the picture there stand for

OpenStudy (anonymous):

10

OpenStudy (anonymous):

100

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} (3\times 10^2)+(8\times 10^1)+(1\times 10^0)\\ \quad \\ \quad \\ \begin{array}{llll} 10^0 & = & 1\\ 10^1 & = & 10\\ 10^2 & = & 100\\ 10^3 & = & 1000\\ 10^4 & = & 10000\\ ... \end{array}\\ \textit{as many zeros as the exponent shows}\\ (3\times 10^2)\implies 3 \times 100 = 300\\ (8\times 10^1)\implies 8 \times 10 = 80\\ (1\times 10^0)\implies 1 \times 1 = 1\\ \quad \\ \quad \\ (3\times 10^2)+(8\times 10^1)+(1\times 10^0) = 300+80+1 \end{array}\)

OpenStudy (jdoe0001):

so 100 ok... can you give me say.... another ending in "0" <- zero

OpenStudy (anonymous):

10\[10^2\]

OpenStudy (anonymous):

Looking at the example from earlier post, I think I get what they want you to practice. Remember that \[1=1*10^{0}.\]. Here is the middle term: \[-514= -1 * ( 5*10^{2} + 1*10^{1} + 4*10^{0} ) = -5*10^{2} - 1*10^{1} - 4*10^{0} \] Then group exponents together and simplify. Example: \[8*10^{2} - 5*10^{2} = (8-5)*10^{2} = 3*10^{2} ==> 800 -500=300\] Does this help? Putting the whole break down and steps into the equation writer would take a while...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so for instance this would be Simplify the expansion https://angel.spcollege.edu/AngelUploads/QuestionData/03263c60-f44c-4b6d-934d-07422994703e/221132662322HR573221.png# {018270b7-a112-45a1-8477-f874d215b8a5}

OpenStudy (anonymous):

10, 795

OpenStudy (jdoe0001):

hmm, ... ok, let's use 10 so we have 100 and 10 the difference is 379 100 + 10 = 110 379 - 110 = 269 so our numbers will be 100 + 10 + 269 \(\bf 100 + 10 + 269\\ \quad \\ \quad \\ 100 = 10^2 \implies 1 \times 10^2\\ 10 = 10^1 \implies 1 \times 10^1\\ 269 = 269 \implies 269 \times 10^0\\ \quad \\ \quad \\ (1 \times 10^2)+(1 \times 10^1)+(269 \times 10^0)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Not quite. We count in base-10. Think of the exponent as the number of zeros that follow the multiplier. your example is 1000 + 700 + 90 + 5 = 1795.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I am lost now

OpenStudy (anonymous):

Sorry... simpler example. Maybe I can get a simpler one in equation format. Give me two three-digit number to subtract in exponential format.

OpenStudy (anonymous):

514-345

OpenStudy (anonymous):

Ah! Good example... give me a minute.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[514-345 => 5*10^2 + 1*10^1 + 4*10^0 - ( 3*10^2 + 4*10^1 + 5*10^0 ) \]

OpenStudy (anonymous):

Working on next transform. (That's the most the equation writer could handle).

OpenStudy (anonymous):

\[5*10^2 +1*10^1+4*10^0 - 3*10^2 - 4*10^1 - 5*10^0\]

OpenStudy (anonymous):

Combine terms by exponent value: \[(5-3)*10^2 + (1-4)*10^1 + (4-5)*10^0\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no that seems easier

OpenStudy (anonymous):

\[2*10^2 - 3*10^1 - 1*10^0\]

OpenStudy (anonymous):

Looking at this now: You can see 200-30-1 = 169 , which is the answer you get from a calculator. Do you need to complete the exercise just with exponents? That is a little more work.

OpenStudy (anonymous):

yes, I know that part

OpenStudy (anonymous):

OK. Did this help?

OpenStudy (anonymous):

Here is the final exponent reduction "just for fun" : \[(2 - .3 - .01)*10^2 = 1.69*10^2 = 169\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

since you explained it , it seems much easier now

OpenStudy (anonymous):

I was lost from the start

OpenStudy (anonymous):

WooHoo!!! I helped someone! ;-) I do these things at work and just get puzzled looks. Work on understanding the details of the last transformation. It's the key to a lot of engineering and chemistry calculations.

OpenStudy (anonymous):

yes :)))

OpenStudy (anonymous):

I know i think i have it down, I wish I could understand my anatomy classes like i understand this